DETECTION OF SiOH (X^{2}A′) BY FOURIER TRANSFORM MICROWAVE SPECTROSCOPY

FILIPPO TAMASSIA, Dipartimento di Chimica Fisica ed Inorganica, Universitá di Bologna, V.le Risorgimento 4, I-40136 Bologna, Italy; \underline{MICHAEL C. MCCARTHY} and PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138.

The rotational spectrum of the free radical SiOH in its X^{2}A′ ground electronic state has been observed by Fourier transform microwave spectroscopy. The 1_{0,1} − 0_{0,0} transition has been detected for the main species ^{28}Si^{16}OH and for the rare isotopic species ^{30}Si^{16}OH, Si^{18}OH, and SiOD. The same transition has also been observed in the excited vibrational states v_{2} = 1 and v_{3} = 1 for the most abundant species. From this data, precise spectroscopic constants have been derived for this plausible astronomical molecule, as has an experimental structure (r_{0}). In addition, the 1_{0,1} −0_{0,0} transition of HNSi has been detected at high spectral resolution by the same technique, yielding a precise value for the nitrogen quadrupole coupling constant.